OpenStudy (anonymous):

How to evaluate 16^-1/4? (Rational exponents)

OpenStudy (anonymous):



OpenStudy (anonymous):

so fraction exponents are like roots so \[16^-{\frac{ 1 }{ 4 }}\] is the same as \[\sqrt[-4]{16}\] which is \[\frac{ 1 }{ \sqrt[4]{16} }\]

OpenStudy (anonymous):

or \[\sqrt[4]{16} = 2 (2 \times its self 4 \times is 16)\]

OpenStudy (anonymous):

the answers is \[\frac{ 1 }{2 }\]

OpenStudy (anonymous):

@TAKEBACKMATH ohhh! okay thank you :)